7-81. The beam consists of three segments pin connected at B and E.Draw the shear and moment diagrams for the beam. 9 kN/m IDShear and Bending Moment Problem 7.81 Engineering Mechanics Statics
Answer #1

Oues) gKN/m A 4.5m72m2m 2me4m E RF RA x4.5x9tox g RA tReRD R PAtRe tRDt Re = 110.2 5 KN Me) et 4.5PA 4-5x9 x4.5 3 RA 6.75 KNManiget) 2 Rc t 4 RD t 10x18 = iDx9 x 2Re 4RD 2 0 -ct Ret2 RD=135 Ret RD= 10.25-675+18) ReRD 85.5 KN olvinged(i) Rc 36 KN RD=- -13 5+49.5= 36 KN ight -18 KN x2 2 x Mx 6.75- M6.75 MA MBO = 1.7 KN m 6.7S X2,626) Mx=26 6.7Sx6,5-x4.5x9 x35- gx2x) Mc = 4518x6-9x6x3 = -54 KN. m MD - ME D (xmght) 18x2-gx2 18 KN.m Mx-2 36KN 6.75KN 4.5KN 18KN 13.5KN 31.5 KN SFD 18KN m 11-7KN m 4SKN